OpenStudy (anonymous):

Let f(x) = 4x – 32 and g(x) = 4x. Part 1 What is (f ◦ g)(x) ? Part 2 Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(4x)=4(4x)-32=16x-32\] your are on your own with the inglish

jhonyy9 (jhonyy9):

sorry i think ,,english"